OpenStudy (anonymous):

Two cars traveled equal distances in different amounts of time. Car A traveled the distance in 2 h, and Car B traveled the distance in 1.5 h. Car B traveled 15 mph faster than Car A. How fast did Car B travel? (Don't just give the answer. Please tell me how you found the answer.)

OpenStudy (anonymous):

let speed of car A = x mph speed of car B = x+15 mph time for car A = 2 hours time for car B = 1.5 hours Since distance is same distance= speed x time 2x = 1.5 ( x+15) Then solve for x.

OpenStudy (anonymous):

answer is 60 mph